OpenStudy (alt23345):

Help with limits: Determine the value of "a" that would make the following function continuous: h(x)={1. x^3 +a^3 if x<2 2. 4-2x if x is greater than or equal to 2 Answer choices: a. -2 b. -1 c. 0 d. 2 e. cannot be determined I have no clue how to figure this out, and I am pretty sure it is very simple.

OpenStudy (evoker):

You need for the value at 2 to be the same for both pieces.

OpenStudy (evoker):

so you need x^3+a^3=4-2x when x=2

OpenStudy (stephy321):

@Evoler, would the answer be a. -2 because when plugged in, both sides of the function equal to 0?

OpenStudy (stephy321):

@Evoker *

OpenStudy (alt23345):

@stephy321 I got A as my answer as well.